OpenStudy (anonymous):

solve solution using elimination 2x+4y=22 2x-2y=-8

OpenStudy (danjs):

multiply one of the equations through by -1 add them together the X term will cancel out 2x-2x=0

OpenStudy (danjs):

2x+4y=22 2x-2y=-8 multiply second by -1 --------------- 2x + 4y = 22 -2x + 2y = 8 add together -------------- 6y = 30 solve for y, then use that y, and solve for x

OpenStudy (danjs):

see that ok?

OpenStudy (anonymous):

y=5

OpenStudy (danjs):

right, i mean, how you got to that point, the work

OpenStudy (anonymous):

by elimination

OpenStudy (danjs):

right of course, i mean do you understand the process that was carried out in order to 'eliminate' one of the variables?